I can't seem figure out the equation for this geometry word problem:
In the illustration, lines r and s are cut by a third line l to form (angle 1) and (angle 2). When lines r and s are parallel, (angle 1) and (angle 2) are supplementary. If (angle 1) = x + 50°,
(angle 2)= 2x - 20°, and lines r and
and s are parallel, find x.


Answer by Theo(13342)

it looks like what they're telling you is that angles 1 and 2 are supplementary because the lines are parallel.

if the angles are supplementary, then their sum is 180 degrees.

so, you get x + 50 + 2x - 20 = 180

combine like termws to get 3x + 30 = 180

subtract 30 from both sides of the equation to get 3x = 150

divide both sides of the equation by 3 to get x = 50

when x = 50, x + 50 = 100 and 2x - 20 = 80

their sum is 180, satisfying the requirement that they be supplementary.

your solution is that x = 50 degrees.